a bicycle wheel turns with constant angular accceleration of 0.64 rad/s²:

a) how much time does it take to reach an angular velocity of 8 rad/sec starting from rest?

b) how many revolutions does the wheel turn in this time interval?

1
Expert's answer
2022-04-11T12:44:45-0400

(a)

"\omega=\omega_0+\beta t"

"t=\frac{\omega-\omega_0}{\beta}=\frac{8-0}{0.64}=12.5\;\rm s"

(b)

"N=\frac{\theta}{2\pi}=\frac{\omega+\omega_0}{2*2\pi}t"

"N=\frac{8+0}{2*2\pi}*12.5=8"